OpenStudy (anonymous):

If f(x) = x² + 1 and g(x) = 3x + 1, find g(20) + f(6). a.74 b.98 c.420

OpenStudy (matt101):

f(6) = 6^2 + 1 = 37 g(20) = 3(20) + 1 = 61 37 + 61 = 98